OpenStudy (anonymous):

Find the constant term in (x/y -y/x)^12

OpenStudy (anonymous):

Wolfram showed 924

OpenStudy (asnaseer):

The constant term would be when both (x/y) and (y/x) have the same power

OpenStudy (asnaseer):

do you know what power that would be?

OpenStudy (anonymous):

the power would be 0

OpenStudy (asnaseer):

no

OpenStudy (experimentx):

Each term is \[ \binom{12}{r} \left ( \frac xy \right )^r \left ( - \frac yx \right )^{12-r}\]

OpenStudy (experimentx):

wher r=0 to 12

OpenStudy (anonymous):

:0 why?

OpenStudy (asnaseer):

Use the hint given by experimentX

OpenStudy (anonymous):

yes, i agree with experimentX

OpenStudy (asnaseer):

so when will: r = 12 - r

OpenStudy (anonymous):

I must make use of the exponents?

OpenStudy (anonymous):

when r =6?

OpenStudy (asnaseer):

r is the exponent of (x/y) and (12-r) is the exponent of (-y/x) --> when will these be equal? yes 6 - that is correct

OpenStudy (asnaseer):

now use the formula given by experimentX to work out the answer

OpenStudy (anonymous):

oooh...ok, the only thing i dont get is how do you know that r=12-r yields a constant term...

OpenStudy (asnaseer):

because:\[(\frac{x}{y})^6\times(\frac{-y}{x})^6\]will cancel all x and y variables

OpenStudy (experimentx):

a bit more general form \[ \binom{12}{r} (-1)^{12-r }\left ( \frac xy \right )^r \left (\frac yx \right )^{12-r} \]

OpenStudy (anonymous):

wouldnt having any exponent that is the same for both terms cancel x and y too?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

and the only place in the expansion of that equation where both exponents are the same is at r=6

OpenStudy (anonymous):

ooooooh, so would you generally use r= n-r; n is the exponent, to see this?

OpenStudy (asnaseer):

in cases like this yes.

OpenStudy (anonymous):

only in cases with 2 different variables?

OpenStudy (asnaseer):

no - only in case where the product of the two variables cancels all "variable"s out.

OpenStudy (anonymous):

so, its more of something that one must think about further?

OpenStudy (asnaseer):

(a-b)^n if a*b results in a constant term then the above rule applies

OpenStudy (anonymous):

what would be the other cases?

OpenStudy (experimentx):

cases where a and b is not symmetric ..... though they might not always ask you to fing the coefficients where power is 0 they might ask you like what is the coefficint of (x/y) .. etc etc

OpenStudy (anonymous):

would \[x ^{3}((x ^{3}/y^2) + (3y/x^2))^9\] be a similar case?

OpenStudy (anonymous):

I'm thinking not..? it doesnt look symmetrical...

OpenStudy (experimentx):

we are looking for ?? with no x and y's??

OpenStudy (anonymous):

yes, the constant term too

OpenStudy (experimentx):

we are looking for 1/x^3

OpenStudy (anonymous):

why is that? will this make the things cancel out too?

OpenStudy (experimentx):

\[ \left ( \frac{x^3}{y^2} \right)^r \left ( \frac{3y}{x^2} \right)^{9-r} = \frac{1}{x^3} \] find the value of r

OpenStudy (anonymous):

3r+9-r=0?

OpenStudy (experimentx):

Oh sorry ... i meant 3

OpenStudy (anonymous):

im not sure what to do now?

OpenStudy (experimentx):

9-r-2r=0 3r - 2(9-r) = -3 solve for r

OpenStudy (anonymous):

r=15?

OpenStudy (experimentx):

it should be less than 9 and must be positive

OpenStudy (anonymous):

oh wait, my bad, its 3

OpenStudy (experimentx):

Umm ... i guess Ohh ... if you cannot find +ve integer value less than .. the given power then you must conclude no solution exist. I hope you can find it's coefficients then

OpenStudy (anonymous):

i think i give up. Its unlikely that will come up in the test anyways. Thank you for your time and help:) much appreciated! :)

OpenStudy (experimentx):

yw ... it's simple!! r = 3 your answer is \[ \binom{9}{3}(3)^{9-3}\]